# B-cos LM: Efficiently Transforming Pre-trained Language Models for Improved Explainability

### v2

@@ Abstract
Post-hoc explanation methods for black-box models often struggle with faithfulness and human interpretability due to the lack of explainability in current neural architectures. Meanwhile, B-cos networks have been introduced to improve model explainability by proposing an architecture that removes bias terms and promotes input-weight alignment. Although B-cos networks have shown success in building explainable systems, their application has so far been limited to computer vision models and their associated training pipelines. In this work, we introduce B-cos LMs, i.e., B-cos language models (LMs) empowered for natural language processing (NLP) tasks. Our approach directly transforms pre-trained language models into B-cos LMs by combining B-cos conversion and task fine-tuning, improving efficiency compared to previous methods. Our automatic and human evaluation results demonstrate that B-cos LMs produce more faithful and human interpretable explanations than post-hoc methods, while maintaining task performance comparable to conventional fine-tuning. Our in-depth analysis explores how B-cos LMs differ from conventionally fine-tuned models in their learning processes and explanation patterns. Finally, we are also the first to explore the transformation of decoder-only models to B-cos LMs for generation tasks.
@@ 1 Introduction
    We present a first study of applying B-cos LMs to decoder-only models and generation tasks, providing a step towards a broader application of B-cos LMs in the era of LLMs.
@@ 3 Methodology
In this section, we outline the architecture and training process of B-cos LMs and how their design ensures faithful and human interpretable explanations. We first introduce B-cos networks (§ 3.1 and § 3.2) and then describe how we transform PLMs to task-specific B-cos LMs (§ 3.3). Finally, we demonstrate how to generate explanations from B-cos LMs (§ 3.4). Notations used in the work are detailed in Appendix B.
@@ 3.1 B-cos Networks
B-cos$\mathbf{(x;w)}$ can be seen as a linear transformation of $\mathbf{x}$ with the dynamic linear weight $\mathbf{w(x)=|\text{cos}(x,\hat{w})|^{\text{B-1}}\times\hat{w}}$. The absence of $\mathbf{b(x)}$ ensures the completeness of summary $\mathbf{w(x)}$. We demonstrate that this completeness extends to an entire network composed of bias-free, dynamic linear modules in § 3.4. Moreover, since the B-cos module output is bounded by $\mathbf{\|x\|}$, the weight $\mathbf{w}$ must align closely with task-relevant patterns to achieve a high cosine similarity and strong activation, especially under additional alignment pressure (B>1). This drives the model to assign greater weight to the most relevant features when optimizing target output probabilities, promoting the learning of representative patterns during training. Consequently, during explanation generation, task-relevant features $\mathbf{x_{i}}$ receive higher attribution $\mathbf{W(x_{i})x_{i}}$ due to stronger alignment, while irrelevant features receive lower attribution, suppressed by weaker alignment and the exponential scaling. For a more detailed discussion of how the B-cos transformation enhances faithfulness and human interpretability, see Böhle et al. (2022; 2024).
@@
3.2 Dynamic Linear Representation of Model Components
B-cos Layers
[…]
Non-linear Activation Functions
Attention Blocks
$\text{MSA}\mathbf{(X)=U[W}_{1}\mathbf{(X)X,...,W}_{H}\mathbf{(X)X)]}$ (4)
Normalization Layers
[…]
@@ 3.4 Computing B-cos Explanations
Once trained, the B-cos LM can generate explanations that faithfully summarize its decision-making process during inference. As all components are dynamic linear with no bias terms (cf. § 3.2), the entire model computation can be expressed as a sequence of matrix multiplications, which can be completely summarized as a single dynamic linear function:
@@ 4 Experiments
We evaluate the task performance of B-cos LMs and faithfulness of B-cos explanations with automatic evaluation across various tasks and PLMs. In addition, we conduct a human evaluation study to compare the human interpretability of B-cos explanations. §4.14.3 describe our automatic evaluation setup, results, as well as human evaluation study, respectively. §4.4 provides a qualitative analysis. Finally, we study the computational efficiency in §4.5 and conduct an ablation study in §4.6. More details on the experimental setup and baseline methods are provided in Appendix C.
@@ 4.1 Experimental Setup
To avoid arbitrary choices of $k$, we compute Comp and Suff for multiple values ($k=10,20,...,90$) and summarize them using the Area Over the Perturbation Curve (AOPC, DeYoung et al., 2020).
@@
Compared to perturbation-based metrics, SeqPG does not rely on perturbations and thus avoids the potential distortions introduced by token masking. When constructing SeqPG examples, we truncate each segment to a fixed length and randomize segment order to control for length and position effects. We generate synthetic examples using correctly and most confidently classified test instances. SeqPG can be seen as a standardized version of hybrid document evaluation (Poerner et al., 2018). We provide an example of SeqPG in Figure 7 and more details in Appendix D.
@@ 4.2 Automatic Evaluation Results
Figure 2 shows the accuracy of conventionally fine-tuned, Saloss and B-cos BERT across three datasets (we provide results for DistilBERT and RoBERTa in Appendix E). On AG News and HateXplain, B-cos LMs performs on par with conventional models, with only a minor drop ($\sim$1%) in accuracy. They also outperform Saloss models on these datasets. Only for IMDB, we find a slightly larger drop of 3.06% compared to conventional BERT, though the performance remains strong overall.
@@
Table 2 shows the faithfulness scores for post-hoc explanation methods on conventionally fine-tuned and Saloss BERT models, as well as B-cos explanations from B-cos BERT. The results show that B-cos explanations are consistently and substantially more faithful than post-hoc methods across all models and datasets. On average, B-cos explanations outperform the strongest post-hoc methods on conventional models by 14.63 points in Comp and achieve negative Suff scores, indicating that the identified important tokens alone enable even more confident predictions. B-cos also shows significant gains in SeqPG. While Saloss improves faithfulness for some post-hoc methods over conventional models, it still underperforms compared to B-cos LMs by a large margin. Similar trends are observed for DistilBERT and RoBERTa (Appendix F) as well, further strengthening our findings. Although we do not include rationale-based models in the main experiments because they typically require additional supervision, a supplementary comparison in Appendix G shows that B-cos BERT still outperforms a rationale-based model on HateXplain.
@@ 4.3 Human Evaluation
Following the practice in Enguehard (2023) and Yue et al. (2022), we randomly select 50 instances, respectively, from AG News and HateXplain where the B-cos and conventional models make the same prediction. Five annotators then rate the explanations in terms of human interpretability (how well they understand them) and human agreement (how much they agree with them) on a scale of 1-5. Before the annotation began, the annotators were provided with a clear description of the evaluation task, metrics, and rating scales. They were also shown example annotations along with the reasoning behind each rating to help them better understand the evaluation criteria. Further details on the evaluation criteria, rating scales, annotator instructions, and example annotations can be found in Appendix H.
@@ 4.4 Qualitative Analysis
Figure 4 provides an example of B-cos and other (post-hoc) explanations. It can be seen that the B-cos explanation highlights important tokens correctly with little focus on irrelevant ones. In contrast, ShapSampl attributes the highest importance to the [SEP] token and provides only little useful information. Meanwhile, DecompX extracts a significant amount of irrelevant information. Overall, we find that the B-cos explanation is more interpretable to humans by providing clearer and more relevant attributions compared to the post-hoc explanations.
Figure 4: Examples of B-cos explanations (B-cos BERT) as well as ShapSampl and DecompX explanations (conv. BERT) from AG News. Green (red) indicates the positive (negative) impact of tokens on the prediction. The B-cos explanation highlights only relevant tokens and is more interpretable to humans (cf. Appendix I for more examples).
4.5 Explanation Efficiency
@@ 4.6 Ablation Study
Method Time (s) Memory (GB)
ShapSampl 37.22 21.95
LIME 6.82 21.96
SIG 67.46 29.09
DecompX 0.76 48.38
B-cos 0.08 2.82
@@
In addition to ablations of model design and training components, we also evaluate alternative explanation methods. Replacing the dynamic linear weights $\mathbf{W(x)}$ with gradients (equivalent to IxG) yields less faithful explanations on B-cos LMs. Besides, directly extracting B-cos-like explanations, $\mathbf{W(x)x}$, from a conventional model results in worse faithfulness compared to those from B-cos LMs. [^3] Extracting $\mathbf{W(x)x}$ from conventional models follows the same approach as in B-cos LMs (cf. § 3.2), except that in standard linear transformations, the dynamic linear weight is replaced by the fixed weight matrix $\mathbf{W}$.
@@ 5 Comparison of B-cosification Setups
We evaluate these setups against the B-cosification approach used in our main experiments (B-cos LM) and compare task performance, faithfulness, and training efficiency (cf. Appendix C for B-cos pre-training details). Additionally, we report results for conventional fine-tuning (Conv. LM) and training a randomly initialized B-cos LM (B-cos from scratch). Experiments are conducted on IMDB with B=1.25 for B-cos models, with results averaged over three runs.
@@
Table 5 shows that B-cos LM requires fewer training steps to reach optimal validation performance than conventional fine-tuning. Training B-cos LM from scratch results in worse task performance and faithfulness, emphasizing the importance of good parameter initialization. Among the two setups that separate task fine-tuning and B-cos conversion, Task then B-cos achieves results similar to B-cos LM but requires more total training steps. B-cos then task initially performs worse under the same training budget. However, with additional pre-training epochs, it surpasses other B-cosification setups in both task performance and faithfulness. Overall, we find that combining task fine-tuning and B-cos conversion is the most efficient approach. However, with sufficient pre-training, B-cos then task can produce more performant and explainable models.
@@ 6 Impact of B-cosification and B Values
For a deeper understanding of how B-cosification and alignment pressure parameter B affect model performance and behavior, we compare conventional and B-cos BERT trained on HateXplain across different B values. We also provide an empirical analysis of the impact of B on input-weight alignment in Appendix J.
@@
Figure 5 shows the effects of varying B on the task performance and explanation faithfulness. Classification accuracy initially improves slightly as B increases from 1 to 1.25, benefiting from the extra non-linearity introduced by B>1. However, beyond this point, accuracy declines as higher alignment pressure reduces model flexibility. A similar trend is observed for Comp, which peaks around B=1.5 before decreasing. This differs from previous findings in vision models (Böhle et al., 2022), which we attribute to the high sparsity of explanations at larger B values. As alignment pressure increases, fewer tokens receive attribution scores that are not close to zero, leading to poor token importance calibration and lower Comp scores. The effects of B on other metrics are similar and can be found in Appendix K.
Explanation Entropy
Figure 5 also reveals a negative correlation between explanation entropy and B, indicating that higher alignment pressure leads to sparser explanations. This aligns with our expectations: a larger B amplifies the differences between dimensions in $\mathbf{|\text{cos}(x,\hat{W})|^{\text{B-1}}}$ of B-cos layers (Equation 2) and the dynamic linear weight assigns more distinct attributions to input features. As a result, explanations become more concentrated, where only a few tokens receive high attributions, while most remain close to zero (cf. Appendix L for an example).
@@
We also find that B-cosification, particularly with large B, amplifies the reliance on spurious correlations. For example, the prediction positive rate for examples with the word “black” rises from 49.02% in the test set and 52.94% in the conventional model to 59.80%, 56.86%, and 73.53% in B-cos LMs with B=1, 1.5, and, 2.5, respectively (we provide an example in Appendix M). However, the faithfulness and interpretability of B-cos explanations facilitate the identification of spurious correlations and can effectively guide models toward reducing them (Rao et al., 2023). We leave the exploration of B-cos LMs for bias detection and mitigation to future work.
7 B-cosifying Decoder-Only Models for Generation Tasks
LLMs are increasingly used as general-purpose assistants, with most based on decoder-only architectures (Zhao et al., 2023; Minaee et al., 2024). While our primary focus is on classification tasks using encoder-only models, we also extend B-cosification to decoder-only models for generation tasks to demonstrate the broader applicability of B-cos LMs. Specifically, we apply B-cosification to two decoder-only models, GPT-2 small (Radford et al., 2019, referred to as GPT-2 afterwards) and Llama-3.2-1B (Dubey et al., 2024, referred to as Llama-3.2 afterwards), and evaluate their language modeling performance and explanation quality on two generation tasks. For more details on the datasets, experimental setup and baseline models, see Appendix C.
@@ 7 B-cosifying Decoder-Only Models for Generation Tasks
Table 6: Language ability results for vanilla and B-cos decoder-only models. Probability Gap scores on BLiMP are averaged over nine subsets. Scores where B-cos LM outperforms their vanilla counterparts are in bold. B-cos LMs show language modeling ability comparable to vanilla models. Results for each subset can be found in Table 13 in Appendix N.
B-cosification Setup
Datasets
For explanation evaluation, we use the BLiMP dataset (Warstadt et al., 2020) to assess explanations for linguistic phenomena, and the Indirect Object Identification (IOI) dataset (Brian Muhia, 2022) to test models’ reasoning about object identification. Following Ferrando et al. (2023), we use nine subsets of BLiMP. Each example in both datasets consists of a sentence prefix followed by a target and a foil next-word prediction, differing in whether they align with the phenomenon or ability of interest. Ground truth evidence is provided to support either grammatical correctness or correct object identification. Examples of these datasets can be found in Table 8 in Appendix C.

### v3

@@ Abstract
Post-hoc explanation methods for black-box models often struggle with faithfulness and human interpretability due to the lack of explainability in current neural architectures. Meanwhile, B-cos networks have been introduced to improve model explainability by proposing an architecture that removes bias terms and promotes input-weight alignment. Although B-cos networks have shown success in building explainable systems, their application has so far been limited to computer vision models and their associated training pipelines. In this work, we introduce B-cos LMs, i.e., B-cos language models (LMs) empowered for natural language processing (NLP) tasks. Our approach directly transforms pre-trained language models into B-cos LMs by combining B-cos conversion and task fine-tuning, improving efficiency compared to previous methods. Our automatic and human evaluation results demonstrate that B-cos LMs produce more faithful and human interpretable explanations than post-hoc methods, while maintaining task performance comparable to conventional fine-tuning. Our in-depth analysis explores how B-cos LMs differ from conventionally fine-tuned models in their learning processes and explanation patterns. Finally, we present a first exploration of transforming decoder-only models to B-cos LMs for generation tasks.
@@ 1 Introduction
    We are also the first to explore the transformation of decoder-only models to B-cos LMs for generation tasks, providing a step towards a broader application of B-cos LMs in the era of LLMs.
@@ 3 Methodology
In this section, we outline the architecture and training process of B-cos LMs and how their design ensures faithful and human interpretable explanations. We first introduce B-cos networks (§ 3.1) and then describe how we transform PLMs to task-specific B-cos LMs (§ 3.2). Finally, we demonstrate how to generate explanations from B-cos LMs (§ 3.3). Notations used in the work are detailed in Appendix B.
@@ 3.1 B-cos Networks
[…]
@@ 3.3 Computing B-cos Explanations
Once trained, the B-cos LM can generate explanations that faithfully summarize its decision-making process during inference. As all components are dynamic linear with no bias terms (cf. Appendix C), the entire model computation can be expressed as a sequence of matrix multiplications, which can be completely summarized as a single dynamic linear function:
@@ 4 Experiments
We evaluate the task performance of B-cos LMs and faithfulness of B-cos explanations with automatic evaluation across various tasks and PLMs. In addition, we conduct a human evaluation study to compare the human interpretability of B-cos explanations. §4.14.3 describe our automatic evaluation setup, results, as well as human evaluation study, respectively. §4.4 provides a qualitative analysis. Finally, we conduct an ablation study in §4.5. More details on the experimental setup and baseline methods are provided in Appendix D and a comparison of computational efficiency is provided in Appendix E.
@@ 4.1 Experimental Setup
To avoid arbitrary choices of $k$, we compute Comp and Suff for multiple values (k=10,20,,90k=10,20,...,90) and summarize them using the Area Over the Perturbation Curve (AOPC, DeYoung et al., 2020).
@@
Compared to perturbation-based metrics, SeqPG does not rely on perturbations and thus avoids the potential distortions introduced by token masking. When constructing SeqPG examples, we truncate each segment to a fixed length and randomize segment order to control for length and position effects. We generate synthetic examples using correctly and most confidently classified test instances. SeqPG can be seen as a standardized version of hybrid document evaluation (Poerner et al., 2018). We provide an example of SeqPG in Figure 7 and more details in Appendix F.
@@ 4.2 Automatic Evaluation Results
Figure 2 shows the accuracy of conventionally fine-tuned, Saloss and B-cos BERT across three datasets (we provide results for DistilBERT and RoBERTa in Appendix G). On AG News and HateXplain, B-cos LMs performs on par with conventional models, with only a minor drop ($\sim$1%) in accuracy. They also outperform Saloss models on these datasets. Only for IMDB, we find a slightly larger drop of 3.06% compared to conventional BERT, though the performance remains strong overall.
@@
Table 2 shows the faithfulness scores for post-hoc explanation methods on conventionally fine-tuned and Saloss BERT models, as well as B-cos explanations from B-cos BERT. The results show that B-cos explanations are consistently and substantially more faithful than post-hoc methods across all models and datasets. On average, B-cos explanations outperform the strongest post-hoc methods on conventional models by 14.63 points in Comp and achieve negative Suff scores, indicating that the identified important tokens alone enable even more confident predictions. B-cos also shows significant gains in SeqPG. While Saloss improves faithfulness for some post-hoc methods over conventional models, it still underperforms compared to B-cos LMs by a large margin. Similar trends are observed for DistilBERT and RoBERTa (Appendix H) as well, further strengthening our findings. Although we do not include rationale-based models in the main experiments because they typically require additional supervision, a supplementary comparison in Appendix I shows that B-cos BERT still outperforms a rationale-based model on HateXplain.
@@ 4.3 Human Evaluation
Following the practice in Enguehard (2023) and Yue et al. (2022), we randomly select 50 instances, respectively, from AG News and HateXplain where the B-cos and conventional models make the same prediction. Five annotators then rate the explanations in terms of human interpretability (how well they understand them) and human agreement (how much they agree with them) on a scale of 1-5. Before the annotation began, the annotators were provided with a clear description of the evaluation task, metrics, and rating scales. They were also shown example annotations along with the reasoning behind each rating to help them better understand the evaluation criteria. Further details on the evaluation criteria, rating scales, annotator instructions, and example annotations can be found in Appendix J.
@@ 4.4 Qualitative Analysis
Figure 4 provides an example of B-cos and other (post-hoc) explanations. It can be seen that the B-cos explanation highlights important tokens correctly with little focus on irrelevant ones. In contrast, ShapSampl attributes the highest importance to the [SEP] token and provides only little useful information. Meanwhile, DecompX extracts a significant amount of irrelevant information. Overall, we find that the B-cos explanation is more interpretable to humans by providing clearer and more relevant attributions compared to the post-hoc explanations. We provide more examples and analyze two undesired explanations from B-cos LMs in Appendix K.
Figure 4: Examples of B-cos explanations (B-cos BERT) as well as ShapSampl and DecompX explanations (conv. BERT) from AG News. Green (red) indicates the positive (negative) impact of tokens on the prediction. The B-cos explanation highlights only relevant tokens and is more interpretable to humans.
@@ 4.5 Ablation Study
In addition to ablations of model design and training components, we also evaluate alternative explanation methods. Replacing the dynamic linear weights $\mathbf{W(x)}$ with gradients (equivalent to IxG) yields less faithful explanations on B-cos LMs. Besides, directly extracting B-cos-like explanations, $\mathbf{W(x)x}$, from a conventional model results in worse faithfulness compared to those from B-cos LMs.
@@ 5 Comparison of B-cosification Setups
We evaluate these setups against the B-cosification approach used in our main experiments (B-cos LM) and compare task performance, faithfulness, and training efficiency (cf. Appendix D for B-cos pre-training details). Additionally, we report results for conventional fine-tuning (Conv. LM) and training a randomly initialized B-cos LM (B-cos from scratch). Experiments are conducted on IMDB with B=1.25 for B-cos models, with results averaged over three runs.
Table 4 shows that B-cos LM requires fewer training steps to reach optimal validation performance than conventional fine-tuning. Training B-cos LM from scratch results in worse accuracy and faithfulness, emphasizing the importance of good parameter initialization. Among the two setups that separate task fine-tuning and B-cos conversion, Task then B-cos achieves results similar to B-cos LM but requires more total training steps. B-cos then task initially performs worse under the same training budget. However, with additional pre-training epochs, it surpasses other B-cosification setups in both task performance and faithfulness. Overall, we find that combining task fine-tuning and B-cos conversion is the most efficient approach. However, with sufficient pre-training, B-cos then task can produce more performant and explainable models.
@@ 6 Impact of B-cosification and B Values
For a deeper understanding of how B-cosification and alignment pressure parameter B affect model performance and behavior, we compare conventional and B-cos BERT trained on HateXplain across different B values. We also provide an empirical analysis of the impact of B on input-weight alignment in Appendix L.
@@
Figure 5 shows the effects of varying B on the task performance and explanation faithfulness. Classification accuracy initially improves slightly as B increases from 1 to 1.25, benefiting from the extra non-linearity introduced by B>1. However, beyond this point, accuracy declines as higher alignment pressure reduces model flexibility. A similar trend is observed for Comp, which peaks around B=1.5 before decreasing. This differs from previous findings in vision models (Böhle et al., 2022), which we attribute to the high sparsity of explanations at larger B values. As alignment pressure increases, fewer tokens receive attribution scores that are not close to zero, leading to poor token importance calibration and lower Comp scores. The effects of B on other metrics are similar and can be found in Appendix M.
Explanation Entropy
@@
Figure 5 also reveals a negative correlation between explanation entropy and B, indicating that higher alignment pressure leads to sparser explanations. This aligns with our expectations: a larger B amplifies the differences between dimensions in $\mathbf{|\text{cos}(x,\hat{W})|^{\text{B-1}}}$ of B-cos layers (Equation 2) and the dynamic linear weight assigns more distinct attributions to input features. As a result, explanations become more concentrated, where only a few tokens receive high attributions, while most remain close to zero (cf. Appendix N for an example).
@@
We also find that B-cosification, particularly with large B, amplifies reliance on spurious correlations. For example, the prediction positive rate for examples with the word “black” rises from 49.02% in the test set and 52.94% in the conventional model to 59.80%, 56.86%, and 73.53% in B-cos LMs with B=1, 1.5, and, 2.5, respectively (we provide an example in Appendix O). However, the faithfulness and interpretability of B-cos explanations facilitate the detection of spurious correlations and can effectively guide models toward reducing them (Rao et al., 2023). We leave the exploration of B-cos LMs for bias detection and mitigation to future work.
7 B-cosifying Decoder-Only Models for Generation Tasks
LLMs are increasingly used as general-purpose assistants, with most based on decoder-only architectures (Zhao et al., 2023; Minaee et al., 2024). While our primary focus is on classification tasks using encoder-only models, we also extend B-cosification to decoder-only models for generation tasks to demonstrate the broader applicability of B-cos LMs. Specifically, we apply B-cosification to two decoder-only models, GPT-2 small (Radford et al., 2019, referred to as GPT-2 afterwards) and Llama-3.2-1B (Dubey et al., 2024, referred to as Llama-3.2 afterwards), and evaluate their language modeling performance and explanation quality on two generation tasks. For more details on the datasets, experimental setup and baseline models, see Appendix D.
@@ 7 B-cosifying Decoder-Only Models for Generation Tasks
B-cosification Setup
Datasets
For explanation evaluation, we use the BLiMP dataset (Warstadt et al., 2020) to assess explanations for linguistic phenomena, and the Indirect Object Identification (IOI) dataset (Brian Muhia, 2022) to test models’ reasoning about object identification. Following Ferrando et al. (2023), we use nine subsets of BLiMP. Each example in both datasets consists of a sentence prefix followed by a target and a foil next-word prediction, differing in whether they align with the phenomenon or ability of interest. Ground truth evidence is provided to support either grammatical correctness or correct object identification. Examples of these datasets can be found in Table 7 in Appendix D.
@@
Table 5: Language ability results for vanilla and B-cos decoder-only models. Scores where B-cos LM outperforms their vanilla counterparts are in bold. B-cos LMs show language modeling ability comparable to vanilla models. Results for each subset can be found in Table 13 in Appendix P.
@@
Results
8 Practical Guidance for Training B-cos LMs
Based on our experiments and analyses, we provide the following guidance on configuring B-cos LMs:
B-cosification Setup
For efficient B-cosification of encoder-only models, we recommend combining B-cos conversion and task fine-tuning. However, if resources permit, an additional B-cos pre-training stage can further enhance both task performance and explanation faithfulness. For decoder-only models, B-cosification can be applied to pre-trained models with unsupervised training data, enabling their conversion into B-cos LMs with improved explainability.
Choice of B value
Several factors influence the selection of an appropriate B value:
(1) Model capacity and explanation sparsity: Excessively large B values can reduce model capacity and produce overly sparse explanations. We therefore recommend setting B within the range of 1–2.
(2) Task complexity and language diversity: For complex tasks with varied language use (e.g., online forum data with diverse language styles), smaller B values are preferable, as they preserve model flexibility and capture more useful features. For other tasks, larger B values can improve explainability.
(3) Bias considerations: Be mindful that B-cosification may amplify biases that the model learns from biased training data. When this risk is present, consider choosing smaller B values and applying bias mitigation methods if necessary.